OpenStudy (anonymous):

Choose the correct simplification of the expression (3x)4. Need help understanding this.

OpenStudy (istim):

Is this multiple choice? What course/grade level is this?

OpenStudy (anonymous):

9th algebra 2

OpenStudy (anonymous):

and yes it is

OpenStudy (istim):

Ok, so what does numbers besides brackets usually represent (HINT- BEDMAS or PEDMAS)

OpenStudy (anonymous):

81x4 12x5 81x5 12x4

OpenStudy (istim):

is x a variable here? OR is that a multiplication sign?

OpenStudy (anonymous):

x variable

OpenStudy (johnweldon1993):

Should we assume this is \[(3x)^4\] ?

OpenStudy (anonymous):

yes

OpenStudy (istim):

Oh. That makes it so much more understandable.

OpenStudy (anonymous):

Sorry I'm just really tired and my mother not letting me get off the laptop till I finsh 12 assignments..

OpenStudy (istim):

Understandable. I think John has some help for you.

OpenStudy (johnweldon1993):

\[(3x)^4 = 3^4 \times x^4\] Since this exponent is separated from the number and variable via the parenthesis...you apply it to each one inside the parenthesis... So what is 3^4?

OpenStudy (anonymous):

81?

OpenStudy (johnweldon1993):

Correct....and what is 81 times x^4?

OpenStudy (anonymous):

81x^4?

OpenStudy (johnweldon1993):

Right...which looks a lot like answer choice...?

OpenStudy (anonymous):

Oh that make sense. Thank you

OpenStudy (anonymous):

Do you think you can help me with another one?

OpenStudy (johnweldon1993):

Right... And remember this can also have been written as \[(3x)^4 = (3x)(3x)(3x)(3x)\] 3x times 3x = 9x^2 so... \[(9x^2)(3x)(3x)\] 9x^2 times 3x = 27x^3 so... \[(27x^3)(3x)\] And finally 27x^3 times 3x = 81x^4 \[(81x^4)\]

OpenStudy (johnweldon1993):

Thanks for the medal by the way @IsTim and sure @BlackLai if I can't then someone else I'm sure will!

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Choose the correct simplification of the expression b5 • b4. b b^9 b^20 b^−1

OpenStudy (johnweldon1993):

Its important to know...that when you are multiplying exponents with the same base...you add the exponents and leave the base alone \[b^5 \cdot b^4 = b^{5 + 4}\]

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

I think I understand abit.

OpenStudy (johnweldon1993):

Yeah just remember some rules \[\large{x^a} \cdot {x ^b} = x^{a + b}\] \[\large \frac{ x^a }{ x^b } = x^{a - b}\] \[\large (x^a)^b = x^{a \times b}\]

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

So for b^5+^4 Do i add add 5+4 then use the exponents to m?ultiply

OpenStudy (anonymous):

multiply*

OpenStudy (johnweldon1993):

wait...what? The question from before? \[b^5 \times b^4\] ?

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

b^5+^4 <----- this what would i have to do get the answer that I need?

OpenStudy (anonymous):

?

OpenStudy (johnweldon1993):

sorry about that...there a mass tagger that keeps calling people to questions lol... you would do b^(5 + 4) and b^9 would be your final answer...just add the exponents...no multiplication here

OpenStudy (anonymous):

Oh okay, you have been really helpful I appericate it.

OpenStudy (anonymous):

So for multiple dat uses the same variable ex: d^3 x d^5

OpenStudy (anonymous):

I would use the same method that you did for my previous question?

OpenStudy (istim):

Yes. but I'm only inferring this from the last few comments.

OpenStudy (johnweldon1993):

Right same method....when the exponents have the same base *as these do..the base is 'd'...you ignore the base...and you add the exponents

OpenStudy (anonymous):

Okay, I believe I understand it now.

OpenStudy (anonymous):

You made it much more clearer for me thank you.

OpenStudy (johnweldon1993):

No problem!

OpenStudy (anonymous):

There's two more i need assist with are you still free to help?

OpenStudy (johnweldon1993):

Yeah sure

OpenStudy (anonymous):

Choose the correct simplification of the expression (a2)3. a5 a8 a a6

OpenStudy (anonymous):

(a^2)^3*

OpenStudy (johnweldon1993):

So remember what I said with these ones \[\large (x^a)^b = x^{a \times b}\]

OpenStudy (anonymous):

Yeah

OpenStudy (johnweldon1993):

So you would use that rule \[\huge(a^2)^3 = a^{2 \times 3} = a^{?}\]

OpenStudy (anonymous):

a^6?

OpenStudy (johnweldon1993):

That is correct!

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Choose the correct simplification of the expression (xyz^2)^4. x^5y^5z^6 x^5y^5z^8 xyz^16 x^4y^4z^8

OpenStudy (johnweldon1993):

Okay now this is just a spin off of the last rule \[\large(xyz^2)^4 = x^4 \times y^4 \times z^{2 \times 4}\] Just remember you apply it to every variable in the equation

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

So how would i get my answer off this one owo

OpenStudy (johnweldon1993):

Well like I showed above...you have x^4 and y^4 and z^(2 times 4) so z^8 Which choice is that?

OpenStudy (anonymous):

Oh I got it now

OpenStudy (anonymous):

Theres this last one if u wish to help me. .?

OpenStudy (johnweldon1993):

Sure why not lol

OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)(2x2y6)

OpenStudy (johnweldon1993):

\[\large(3x^3y^4)(2x^2y^6)\] like that?

OpenStudy (anonymous):

(3x^3y^4)(2x^2y^6)*

OpenStudy (anonymous):

ye

OpenStudy (johnweldon1993):

So we can break this up a bit The numbers first... what is 3 times 2?

OpenStudy (anonymous):

6

OpenStudy (johnweldon1993):

Right... what is x^3 times x^2? *hint* \[\huge x^3 \times x^2 = x^{3 + 2}\]

OpenStudy (anonymous):

X^3+2 = X^5

OpenStudy (johnweldon1993):

Right....and finally what is y^4 times y^6? *hint* \[\huge y^4 \times y^6 = y^{4 + 6}\]

OpenStudy (anonymous):

Y^4+6= Y^10

OpenStudy (johnweldon1993):

Great...so altogether we have \[\large 6x^5y^{10}\]

OpenStudy (anonymous):

So no signs in the middle of those equations?

OpenStudy (johnweldon1993):

None at all...there were none to begin with so there are none to end with

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

In a way how would I explain it in complete sentence?

OpenStudy (johnweldon1993):

Hmm...when you multiply these 2 together....you multiply the coefficients.....add the exponents of the base 'x' and add the exponents of the base 'y' to get 6x^5y^10 Idk something like that I guess lol

OpenStudy (anonymous):

lol That understandable

OpenStudy (anonymous):

Well that all for now thank you again for all your help.

OpenStudy (johnweldon1993):

lol no problem man!